OpenStudy (anonymous):

Two children, who are 224 meters apart, start walking toward each other at the same instant at rates of 1.5 m/sec and 2 m/sec. When will they meet? How far will each have walked?

OpenStudy (anonymous):

2-(1.5+2)t=0

OpenStudy (anonymous):

...what?

OpenStudy (anonymous):

2 - (1.5 + 2) t=0 solve for t

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

at t=0 , the distance between them is 2 at t=.1 2- 1.5(.1) -2.0(.1)=1.65